# GenVC: Self-Supervised Zero-Shot Voice Conversion

### v1

@@ IV Experiments
We conduct a series of experiments to evaluate the effectiveness of our proposed VC approach. Using English speech data, we examine two configurations of the GenVC system: one trained on a smaller dataset and the other on a large-scale dataset. Both subjective and objective evaluations are performed to assess the quality of the converted speech, benchmarking our systems against several state-of-the-art models. Furthermore, we analyze the privacy-preserving capabilities of the proposed systems by measuring the extent to which the original speaker’s style is retained or leaked in the converted speech.
@@ IV-A Dataset
For evaluation, we used a range of test sets commonly used in VC and anonymization studies. Specifically, the CMU Arctic [45] and Emime [46] datasets are employed to construct source-target conversion pairs, enabling the assessment of voice conversion performance. Furthermore, in accordance with the speech anonymization pipeline defined by the VoicePrivacy Challenge 2024 (VPC2024) [47], we use the Librispeech dataset [48] to evaluate linguistic content preservation and privacy protection. In our setup, the pseudo-anonymized voices are sourced from the VoxCeleb1 dataset [12].
@@ IV-B Training and Inference Details
The HiFiGAN vocoder’s generator and discriminator are optimized using the AdamW algorithm with a learning rate of $2\times 10^{-4}$. The HiFiGAN generator contains 3.16M parameters. Additionally, four types of discriminators are employed for vocoder training: multi-scale discriminator, multi-period discriminator, multi-scale STFT discriminator, and multi-scale sub-band constant-Q transform discriminator [51]. Models We trained two models, GenVC-Small and GenVC-Large, for this study. Both models use the same audio tokenizers trained on the LibriTTS dataset but differ primarily in the training data used for Phase 2 and Phase 3.
@@
GenVC-Large is fine-tuned initialized from GenVC-Small using additional data from the CommonVoice-EN and MLS-EN datasets, totaling approximately 45k hours of audio. This training was performed with the Perceiver encoder kept frozen. The language model was fine-tuned with a batch size of 24 for 1.5M steps, while the HiFiGAN vocoder was trained with a batch size of 128 for 220k steps.
@@ V-A Conversion Performance
We evaluate the performance of our proposed systems using conversion pairs constructed from the test set. For comparison, we include three previously established state-of-the-art zero-shot systems—YourTTS, FreeVC, and Neural VC—since their code and model checkpoints are publicly available for inference. Note that these baseline systems were trained on the VCTK dataset and rely on an external ASV system to render speaker identity.
@@ V-A1 Objective Evaluation
The results are presented in Table II. YourTTS achieves the highest speaker similarity (0.885), closely followed by Neural VC (0.881) and FreeVC (0.878). GenVC-Large (0.884) surpasses FreeVC and Neural VC, while GenVC-Small (0.869) remains competitive. Overall, similarity scores are closely aligned. For naturalness, FreeVC leads (UTMOS 2.81), with GenVC-Small (2.66) and GenVC-Large (2.65) slightly outperforming Neural VC (2.60) and significantly exceeding YourTTS. These results highlight GenVC’s balanced performance, achieving competitive speaker similarity and naturalness against leading zero-shot VC systems.
@@ V-A2 Subjective Evaluation
For the subjective evaluation test, we randomly selected 50 pairs of utterances for evaluation. 11 participants were recruited to conduct the listening test, where they rated the naturalness and similarity of both genuine utterances and voice converted utterances. Each participant rated a total of 330 utterances: 250 utterances from five VC systems, 80 genuine audio samples. The stimuli were presented in random order, and detailed instructions are provided in Appendix B.
The naturalness mean opinion score (NMOS) and similarity mean opinion score (SMOS) results are summarized in Table III. According to the results, FreeVC achieves the highest NMOS of 3.92, indicating the best perceived naturalness among the evaluated VC systems. In comparison, GenVC-Small and GenVC-Large achieve lower NMOS scores of 3.46 and 3.32, respectively. However, it is important to note that the quality of training data plays a crucial role in the perceived naturalness. The baseline systems are trained on VCTK, which yields higher NMOS scores compared to the training sets used for GenVC, suggesting that differences in dataset quality may have influenced the results.
@@
     VCTK 4.19 $\pm$ 0.13 -
     Train-Small 4.14 $\pm$ 0.15 -
     Train-Large 4.02 $\pm$ 0.09 -
     YourTTS 2.87 $\pm$ 0.09 2.87 $\pm$ 0.08
     FreeVC 3.92 $\pm$ 0.08 3.48 $\pm$ 0.07
     Neural VC 3.86 $\pm$ 0.08 3.38 $\pm$ 0.07
     GenVC-Small 3.46 $\pm$ 0.09 3.68 $\pm$ 0.08
     GenVC-Large 3.32 $\pm$ 0.09 3.49 $\pm$ 0.08
Our proposed VC approach demonstrates strong performance in speaker similarity. GenVC-Small achieves the highest SMOS of 3.68, surpassing all other systems, including the best-performing baseline, FreeVC, which scores 3.48. However, GenVC-Large does not show further improvement in similarity despite being trained on a larger dataset, though it still outperforms the baseline systems. This may be partially attributed to the quality of the additional training data. Overall, the results suggest that the GenVC models achieve state-of-the-art speaker similarity while maintaining competitive naturalness. Nevertheless, there exists a trade-off: incorporating larger but lower-quality training data can degrade conversion quality.
@@ V-B Anonymization Performance
Although VC can effectively anonymize voices in terms of timbre, many VC-based approaches reveal source speaker information through the prosodic structure of the speech, thereby limiting privacy preservation. Here we adopt the anonymization and evaluation pipeline from the VoicePrivacy Challenge 2024 [47] to assess our models’ anonymization performance in comparison with other VC-based approaches.
The privacy evaluation pipeline follows a standard speaker verification process. In this setup, the verification model is trained on anonymized data, with labels corresponding to the original speaker identities. A successful anonymization system should sufficiently distort and obscure the original speaker’s identity, making it difficult for the verification system to accurately identify the source speaker. During evaluation, pairs of source speech, referred to as trials, from the evaluation dataset are anonymized and treated as both enrollment and test speech. The primary metric for privacy evaluation is the equal error rate (EER), calculated based on similarity scores from these pairs. A lower EER indicates a higher risk of speaker re-identification, while a higher EER reflects better performance in preserving voice privacy. In addition, utility evaluation is performed to measure the preservation of linguistic content. The anonymized utterances are transcribed using a speech recognition system, and the resulting transcripts are compared with the ground-truth content from the source data. The word error rate (WER) is used as the metric, with a lower WER indicating better preservation of the linguistic content.
For this experiment, the verification models are trained on converted utterances from the train-clean-360 set of Librispeech and tested on the development and test sets of Librispeech. Target voices are randomly selected from VoxCeleb1.
As shown in Table IV, other VC-based anonymization approaches present risks of revealing the source speaker’s identity, as their average EERs are close to or below 12%. YourTTS, FreeVC, and Neural VC demonstrate similarly limited privacy preservation in this context. In contrast, the GenVC models show significant improvements over the other systems. Specifically, the GenVC-Small model achieves an EER of 28.84%, while the GenVC-Large model achieves an EER of 26.88%, outperforming the best VC-based approach (YourTTS) by over 15%.
TABLE IV: Anonymization performance of VC-based approaches
@@
YourTTS 11.8 6.09
FreeVC 9.51 2.7
Neural VC 10.12 3.53
GenVC-Small 28.84 6.87
GenVC-Large 26.88 6.72
We note that while Table IV suggests that the improved privacy preservation in the GenVC model may have resulted in part from poorer content preservation, this is likely an artifact of our training procedure, which used significantly shorter speech chunks during training. We experimented with training GenVC-Small using full contexts for SSL feature extraction, which led to improved privacy preservation without degradation in content preservation (3.55% WER). However, this approach is less efficient for large-scale training and requires substantial storage for offline feature extraction. In this paper, we report on models that are sufficiently effective for large-scale on-the-fly training, but we plan to further improve their performance to align with the offline feature extraction approach in future work.
VI Discussions
Our work has several limitations, primarily due to computational constraints. For instance, we did not explore the impact of model size on performance, despite scaling laws playing a critical role in optimizing LMs for large-scale training [54]. Additionally, we did not investigate the influence of acoustic and phonetic token configurations or evaluate alternative SSL models, which could offer valuable insights compared to the ContentVec model used in our experiments.
@@ VI Discussions
A significant challenge in zero-shot VC remains the accurate preservation of speaker timbre—especially for unseen voices—alongside other speaker-specific characteristics such as accent, emotion, prosody, and recording environment [13, 33]. These factors are crucial for both identifying target speakers and anonymizing source voices. Addressing this challenge requires the availability of more diverse training data, including recordings from varied acoustic environments [24]. While such data may not always match the quality of studio-recorded utterances, it could significantly improve the robustness of VC models in real-world applications. We believe that self-supervised learning approaches, such as GenVC, hold promise for enabling scalable training and improving generalization to unseen speakers and conditions.
Future work will focus on addressing several limitations and improving the model’s robustness. Specifically, we aim to optimize GenVC’s streaming capabilities and enhance its robustness for multilingual and cross-lingual VC. Expanding the model’s applicability across diverse linguistic and acoustic environments is an essential step toward developing more practical, adaptable, and inclusive VC systems.
@@ VII Conclusions
This paper introduces GenVC, a self-supervised zero-shot voice conversion system designed to leverage in-the-wild data for training. Compared to other LM-based approaches, it features a simpler structure and eliminates the need for external supervised models. It effectively learns the disentanglement of linguistic content from speaker-specific attributes without relying on supervised models. Our experimental results show that GenVC delivers outstanding voice cloning capability while maintaining naturalness competitive with state-of-the-art zero-shot voice conversion systems. Moreover, GenVC effectively modifies the prosodic characteristics of the source speech, significantly enhancing privacy preservation in anonymization applications.
@@ Acknowledgment
Impact Statement
This paper presents work aimed at advancing the field of Voice Conversion. While our work holds significant promise, it also carries potential societal implications that warrant consideration. GenVC is a powerful voice conversion system capable of transforming source speech into desired voices. While this technology has valuable applications, such as enhancing privacy by anonymizing voices and enabling accessibility for individuals with speech impairments, it is also presents ethical challenges. Specifically, the ability to convincingly replicate voices can be misused to create audio deepfakes, which may be employed for malicious purposes, such as identity theft, fraud, and the spread of misinformation.
To mitigate these risks, we strongly advocate for the responsible and ethical use of voice conversion technologies. Researchers, developers, and users must comply with relevant laws and guidelines, ensuring that these systems are used exclusively for legitimate and beneficial applications. Transparency, informed consent, and robust safeguards should be prioritized to prevent misuse and protect individuals’ rights and privacy.
Furthermore, we emphasize the importance of continued research into developing countermeasures, such as deepfake detection tools, and fostering collaboration between researchers, policymakers, and industry stakeholders to address the ethical implications of voice conversion technology. By proactively addressing these concerns, we can maximize the benefits of systems like GenVC while minimizing their potential for harm. To support responsible use, we are releasing our code and models to the research community.
Acknowledgment
Appendix A Model Architecture
A-A DVAE
Our DVAE encoder and decoder are both made up of a sequence of 1-D convolutional layers. The first two encoder convolutional layers upsample the input dimensionality to the DVAE hidden dimension of $1024$. They are followed by $3$ Resblocks [55] and a final 1-D convolutional layer that projects the representation from the DVAE hidden dimension to the codebook dimension of $512$. Conversely, the decoder is a mirror image of the encoder, consisting the above-mentioned encoder components but in reverse order.
A-B Perceiver Encoder
Our Perceiver encoder module comprises learned latent queries and a cross-attention mechanism. The learned queries attends to all input frames, transforming them into fixed-length representations. The encoder consists of four multi-head attention blocks, each with 8 heads, where each head has an inner dimension of 64.
Appendix B Subjective Evaluation Instructions
B-A Naturalness
In this experiment, please listen to the speech sample and rate their naturalness on a scale from 1 (Bad) to 5 (Excellent), with increments of 0.5. 1, 1.5, 2, 2.5, 3, 3.5, 4, 4.5 and 5 are allowed on the half-point scale. The scale is defined as follows:
  1. 1.
    Bad: Very unnatural speech, completely unrecognizable as human speech.
  2. 2.
    Poor: Noticeably unnatural speech with many artifacts and discontinuities.
  3. 3.
    Fair: Moderately unnatural speech with noticeable artifacts, but not entirely unnatural.
  4. 4.
    Good: Mostly natural speech with only minor artifacts that are noticeable upon careful listening.
  5. 5.
    Excellent: Perfectly natural speech, indistinguishable from human speech.
B-B Similarity
In this experiment, you will listen to pairs of speech samples and rate how similar the second sample sounds to the reference speech in terms of speaker voice, speaking style, and environmental background, regardless of the content (which will differ). Use a scale from 1 (Bad) to 5 (Excellent), with increments of 0.5. The allowed ratings are 1, 1.5, 2, 2.5, 3, 3.5, 4, 4.5, and 5. The scale is defined as follows:
  1. 1.
    Bad: The two samples sound completely different with barely any similarity.
  2. 2.
    Poor: The two samples have some resemblance, but the differences are significant. For example, the speakers may differ in gender or pitch (e.g., a high-pitched female voice vs. a low-pitched male voice). Speaking style and environmental background also differ noticeably.
  3. 3.
    Fair: The two samples sound somewhat similar, but there are still noticeable differences. It’s clear the speakers are the same gender, but their voices are distinctly different, and their speaking styles differ somewhat.
  4. 4.
    Good: The given utterance sounds quite similar to the reference utterance, with only minor differences noticeable upon close listening. The speaker voices are close, and the speaking styles largely match.
  5. 5.
    Excellent: The given utterance sounds identical to the reference utterance, with no perceivable differences. The timbre of the speakers is the same, their speaking styles match perfectly, and the environmental background is similar, including any acoustic noise.

### v2

@@ IV Experiments
We conduct a series of experiments to evaluate the effectiveness of our proposed VC approach. Using English speech data, we examine two configurations of the GenVC system: one trained on a smaller dataset and the other on a large-scale dataset. Both subjective and objective evaluations are performed to assess the quality of the converted speech, benchmarking our systems against baseline models. Additionally, we analyze GenVC’s privacy-preserving capabilities and evaluate speaker information preservation from different modules to validate speech disentanglement.
@@ IV-A Dataset
For evaluation, we use commonly used test sets in VC and anonymization studies. Specifically, the CMU Arctic [49] and Emime [50] datasets are used to construct conversion pairs, enabling the assessment of zero-shot VC performance. Furthermore, in accordance with the speech anonymization pipeline defined by the VoicePrivacy Challenge 2024 (VPC2024) [37], we use the Librispeech dataset [51] to evaluate linguistic content preservation and privacy protection. We also include VoxCeleb datasets [52, 53] for experiments regarding speaker disentanglement.
@@ IV-B Training and Inference Details
The vocoder’s generator and discriminator are optimized using the AdamW algorithm with a learning rate of $2\times 10^{-4}$. The HiFiGAN generator contains 3.16M parameters. Additionally, four types of discriminators are employed for vocoder training: multi-scale discriminator, multi-period discriminator, multi-scale STFT discriminator, and multi-scale sub-band constant-Q transform discriminator [56]. Models We trained two models, GenVC-Small and GenVC-Large, for this study. Both models use the same audio tokenizers trained on the LibriTTS dataset but differ primarily in the training data used for Phase 2 and Phase 3.
@@
GenVC-Large is fine-tuned and initialized from GenVC-Small using additional data from the CommonVoice-EN and MLS-EN datasets. This training was performed with the Perceiver encoder kept frozen. The language model was fine-tuned with a batch size of 24 for 1.5M steps, while the HiFiGAN vocoder was trained with a batch size of 128 for 220k steps.
@@ V-A Conversion Performance
We evaluate the performance of our proposed systems using conversion pairs constructed from the test set. For comparison, we include three leading zero-shot systems—YourTTS, FreeVC, and Neural VC—since their code and pretrained checkpoints are publicly available for inference. Similar approaches, such as LM-VC [16] and StreamVoice [35], are excluded due to the lack of released implementations. Note that the selected baseline systems were trained on the VCTK dataset and rely on an external ASV system to render speaker identity.
@@ V-A1 Objective Evaluation
The results are presented in Table II. YourTTS achieves the highest speaker similarity, closely followed by Neural VC and FreeVC. GenVC-Large surpasses FreeVC and Neural VC, while GenVC-Small remains competitive. Overall, similarity scores are close among all systems. For naturalness, FreeVC achieves the best performance, while GenVC-Small and GenVC-Large slightly outperform Neural VC and significantly exceed YourTTS.
@@ V-A2 Subjective Evaluation
For the subjective evaluation test, we randomly selected 50 pairs of utterances for conversion. A total of 13 participants were recruited to perform a listening test, in which they rated the naturalness and speaker similarity of genuine and voice-converted utterances. Ratings were given on a scale from 1 (Bad) to 5 (Excellent), with 0.5-point increments. Each participant rated a total of 330 utterances: 250 utterances from five VC systems, 80 genuine audio samples. The stimuli are presented in random order, and detailed instructions are provided on the sample webpage.
The naturalness mean opinion score (NMOS) and similarity mean opinion score (SMOS) results are summarized in Table III. We use the Mann-Whitney U test to assess the statistical significance of all possible pairwise comparisons between different VC approaches [59, 60]. Most pairwise differences in MOS are statistically significant (p-value <0.05<0.05), except for the NMOS and SMOS comparisons between NeutralVC and FreeVC, and the SMOS comparison between FreeVC and GenVC-Large.
@@
     VCTK 4.23 $\pm$ 0.12 -
     Train-Small 4.15 $\pm$ 0.14 -
     Train-Large 4.04 $\pm$ 0.08 -
     YourTTS 2.96 $\pm$ 0.09 2.87 $\pm$ 0.07
     FreeVC 3.96 $\pm$ 0.07 3.48 $\pm$ 0.07
     Neural VC 3.89 $\pm$ 0.08 3.39 $\pm$ 0.07
     GenVC-Small 3.57 $\pm$ 0.09 3.72 $\pm$ 0.08
     GenVC-Large 3.43 $\pm$ 0.09 3.50 $\pm$ 0.08
According to the results, FreeVC achieves the highest NMOS of 3.96, indicating the best perceived naturalness among the evaluated VC systems. In comparison, GenVC-Small and GenVC-Large achieve lower NMOS scores of 3.57 and 3.43, respectively. However, the quality of training data might impact perceived naturalness. While the baseline systems were trained on the high-quality VCTK dataset, GenVC was trained on datasets with comparatively lower quality, which may have contributed to the observed differences in NMOS.
Our proposed VC approach demonstrates strong performance in speaker similarity. GenVC-Small achieves the highest SMOS of 3.72, surpassing all other systems, including the best-performing baseline, FreeVC, which scores 3.48. However, GenVC-Large does not show further improvement in similarity despite being trained on a larger dataset, though it still outperforms the baseline systems. This may be partially attributed to the quality of the additional training data. Overall, the results suggest that the GenVC models achieve state-of-the-art speaker similarity while maintaining competitive naturalness. Nevertheless, a key limitation remains: although the model’s self-supervised nature supports large-scale training, incorporating larger but lower-quality datasets can negatively affect conversion quality.
@@ V-B Anonymization Performance
Although VC can effectively anonymize voices in terms of timbre, many VC-based approaches reveal source speaker information through the prosodic structure of the speech, thereby limiting privacy preservation. Here we adopt the anonymization and evaluation pipeline from VPC2024 [37] to assess our models’ anonymization performance in comparison with other VC-based approaches.
The privacy evaluation pipeline follows a standard speaker verification process. In this setup, the verification model is trained on anonymized data labeled with the original speaker identities. A successful anonymization system should sufficiently distort and obscure the original speaker’s identity, making it difficult for the verification system to accurately identify the source speaker. During evaluation, pairs of source speech, referred to as trials, from the evaluation dataset are anonymized and treated as both enrollment and test speech. The primary metric for privacy evaluation is the equal error rate (EER), calculated based on similarity scores from these pairs. A lower EER indicates a higher risk of speaker re-identification, while a higher EER reflects better performance in preserving voice privacy. In addition, utility evaluation is performed to measure the preservation of linguistic content. The anonymized utterances are transcribed using a speech recognition system, and the resulting transcripts are compared with the ground-truth content from the source data. The word error rate (WER) is used as the metric, with a lower WER indicating better preservation of the linguistic content.
For this experiment, the verification models are trained on converted utterances from the train-clean-360 set of Librispeech and tested on the development and test sets of Librispeech. Target voices are randomly selected from EMIME. As shown in Table IV, other VC-based anonymization approaches present risks of revealing the source speaker’s identity, as their average EERs are close to or below 12%. YourTTS, FreeVC, and Neural VC demonstrate similarly limited privacy preservation in this context. In contrast, the GenVC models show significant improvements over the other systems. Specifically, the GenVC-Small model achieves an EER of 29%, while the GenVC-Large model achieves an EER of 27%, outperforming the best VC-based approach (YourTTS) by over 15%.
TABLE IV: Anonymization performance of VC-based approaches.
@@
YourTTS 11.8 6.31
FreeVC 9.1 2.73
Neural VC 10.4 3.50
GenVC-Small 29.0 6.03
GenVC-Large 27.0 5.65
In terms of content preservation, FreeVC performs best among the evaluated VC approaches. The GenVC models achieve a WER of approximately 6%, comparable to that of YourTTS, indicating sufficient preservation of linguistic content.
V-C Speaker-Content Disentanglement
The preceding subjective evaluation and synthesized samples demonstrate that our model effectively converts source speech into target voices while preserving linguistic content. Nevertheless, speech signals inherently encode multiple attributes—such as speaker identity and environmental context—which our model may implicitly learn to disentangle during conversion. In this experiment, we focus specifically on the disentanglement of speaker identity. To assess the extent to which various model components preserve speaker-specific information, we construct speaker verification systems using features extracted from the trained phonetic tokenizer and the Perceiver encoder, treating these modules as fixed feature extractors.
The first system uses phonetic tokens as input features, with an embedding table that maps each token to a continuous vector representation. We employ the ECAPA-TDNN architecture [61] to train a speaker verification model using the VoxCeleb1 and VoxCeleb2 datasets, without any data augmentation. During evaluation, verification scores are computed using cosine similarity. As shown in Table V, this system yields an equal error rate (EER) close to 50%, suggesting that the phonetic token representations retain minimal speaker-specific information. In contrast, System 2 follows the same training and evaluation protocol but utilizes the output of the Perceiver encoder from the GenVC-Small model as input features. This system achieves a markedly lower EER of 17.7%, indicating that the Perceiver encoder retains a significant amount of speaker-related information. It is worth noting that the Perceiver encoder is trained in an unsupervised fashion on the LibriTTS dataset, which contains approximately 2,300 speakers. As a result, it may not serve as an ideal feature extractor for speakers in the VoxCeleb datasets.
TABLE V: The performance of verification models on VoxCeleb1-O.
ID Source Feat. Front-end + Back-end EER(%) $\downarrow$
1 Phonetic-Token ECAPA-TDNN + Cos 48.5
2 Perceiver-Feat. ECAPA-TDNN + Cos 17.7
3 Perceiver-Mean Cos 34.7
4 Perceiver-Mean PLDA 12.2
We further evaluate the speaker-discriminative capability of the Perceiver encoder by applying simple back-end verification methods. System 3 computes the mean of the Perceiver output across time and uses cosine similarity as the scoring metric. It achieves an EER of 34.7% on the VoxCeleb1-O test set. In System 4, we replace cosine scoring with a probabilistic linear discriminant analysis (PLDA) backend trained on the VoxCeleb datasets using the same mean Perceiver features, which significantly improves performance to an EER of 12.2%. These results provide additional evidence that the Perceiver encoder effectively captures speaker-specific information.
Overall, our findings highlight the effectiveness of the proposed model in sufficiently disentangling speaker identity during training: while the content tokens extracted from source speech contain minimal speaker information, the Perceiver encoder captures most of the speaker-related characteristics.
VI Limitations and Future Work
@@
A significant challenge in zero-shot VC remains the accurate cloning of speaker timbre, especially for unseen voices, alongside other speaker-specific characteristics such as accent, emotion, prosody, and recording environment [6, 16]. These factors are essential for both achieving expected target voice conversions and effectively anonymizing source voices. Addressing this challenge requires access to more diverse training data, including recordings from varied acoustic environments [26]. While such data may not match the quality of studio-recorded utterances, it could substantially improve the robustness of VC models in real-world applications. We believe that self-supervised learning approaches, such as GenVC, hold promise for enabling scalable training and enhancing generalization to unseen speakers and conditions.
However, our current study is subject to several limitations, mainly due to computational constraints. For instance, we did not explore the impact of model size on performance, despite scaling laws playing a critical role in optimizing LMs for large-scale training [62]. We did not investigate the influence of acoustic and phonetic token configurations or evaluate alternative SSL models, which could offer valuable insights compared to the ContentVec model used in our experiments. Addressing these gaps will be a key focus of future work. In addition, optimizing GenVC’s streaming capabilities and enhancing its robustness for multilingual and cross-lingual VC can be another interesting future work.
@@ VII Conclusions
This paper introduces GenVC, a self-supervised zero-shot voice conversion system designed to leverage in-the-wild data for training. Unlike other LM-based approaches, GenVC adopts a simpler architecture and does not rely on external supervised models to disentangle linguistic content and speaker-specific attributes. Our experimental results show that GenVC delivers outstanding voice cloning capability while maintaining naturalness competitive with state-of-the-art zero-shot voice conversion systems. Moreover, GenVC effectively modifies the prosodic characteristics of the source speech, significantly enhancing privacy preservation in anonymization applications.